# RoboHanger: Learning Generalizable Robotic Hanger Insertion for Diverse Garments

### v1

@@ III-E Data Collection
We use FEM and implicit time integration with a single-step Newton iteration [67] for physical simulation. Cloth self-collision and cloth-to-rigid-body collision are handled with a cubic energy function. Cloth self-intersection is prevented using Continuous Collision Detection (CCD). The collision force between cloth and rigid bodies is computed through the Signed Distance Function (SDF) of the latter. We do not perform dynamics simulation or collision handling for rigid bodies and articulated bodies; instead, we compute the object’s position and velocity using kinematic relationships. We render depth images with SAPIEN [68] and design the 3D models of clothes and hangers (without texture) heuristically.
Domain randomization
We randomize the Young’s modulus and density of the clothing, as well as the friction coefficient of the garment in the simulation. Additionally, we randomize the camera’s intrinsic and extrinsic parameters to compensate for calibration errors when the system is deployed in the real world.
@@
We use an iterative data collection approach. Specifically, we first design a state-based heuristic policy to collect trajectories, and use this data to train our networks for initialization. The heuristic policy attempts to find an action that meets all of our predefined rules through sampling. These rules are based on geometric information of the current clothing mesh, such as the number of layers of the garment near the grasping point. Next, we use the network policy to collect more data in the simulator, incorporate this online data into the dataset, and iteratively finetune the network.
III-F Whole-body Motion Controller
@@ IV Experiment
IV-A Experiment Setup
Hanger. We find that using parallel grippers makes it difficult to reliably grasp thin hangers. Therefore, we employ custom-designed 3D-printed hangers to facilitate grasping. The primary difference between our hangers and regular ones is that our design includes a section specifically made for secure gripping.
Metrics. We use the success rate of inserting the hanger, $S$, to evaluate the performance of the policy. $S$ can be further divided into: the success rate of inserting the left endpoint, $S_{1}$; and the success rate of inserting the right endpoint after successfully inserting the left endpoint, $S_{2}$. In our policy, the left end is always inserted first, followed by the right end. Therefore, $S=S_{1}\times S_{2}$.
@@ IV-B Comparison with Baselines in Simulation
We first compare our methods with several other baselines in simulation. Among the 144 synthetic clothing meshes, we randomly divided them into 120 training garments and 24 testing garments. The hangers used for training and testing are the ones shown in Figure 4 (a.1) and (a.2).
@@
We present the comparisons in Table I. ”Input” refers to the type of input used by the policy, which can be either visual input (V) or state input (S). ”Output” refers to the type of output generated by the policy, which can either be action primitives (Prim) or joint angles (Joint).
”Heuristic” represents a heuristic algorithm that we manually designed based on the current mesh of the clothing.
”IL” stands for imitation learning, which is also based on action primitives and UNets. We treat the task as a multi-classification task (selecting a pixel in the image) using cross-entropy loss and only retain positive samples for training the UNets.
”SAC” stands for Soft Actor-Critic algorithm [69]. In SAC, the input to the value network consists of both the observation and the action. The action is first transformed into an image where each pixel value represents the inverse of the distance to that action. This action image is then concatenated with the observation and fed into a CNN. The action network is also a CNN that directly outputs $(X_{i},Y_{i})$. Compared to our method, SAC requires learning an additional action network, while our method directly finds the maximum position on the output of the value network.
”ACT” and ”DP” (diffusion policy) are methods proposed by [43] and [70]. We directly use data generated by heuristics to train an end-to-end imitation learning policy. The network’s input consists of the raw depth images from the head camera as well as the masks of the clothing and the hanger.
Our method performs best in simulation and achieves an overall 85.1% success rate. ACT performs excellently when inserting the first endpoint, but due to accumulated errors, it performs poorly when inserting the second endpoint. Similar issues are observed in other end-to-end imitation learning baselines such as diffusion policy. We also observe that SAC performs worst probably due to the additional action network not sufficiently trained.
@@ IV-C Ablation Studies in Simulation
”w/o online data” indicates training the network using only offline data without online data, under the same data quantity. Fine-tuning with online data leads to an improvement in the final success rate. ”w/o neckline mask” indicates training the policy without using the neckline mask. It is evident that the neckline mask also contributes to an increase in the final success rate.
@@ IV-D Comparison with Baselines in Real World
In real-world experiments, we test 8 different garments and Hanger 1, as shown in Figure 4. For a fair comparison, we use the robot to initialize the garment configuration. Additionally, to evaluate our policy’s performance under different initializations, we also report the policy’s success rate with human initialization.
@@
We present the performance of our policy compared to other methods on the real robot in Table III. We include ”IL” (imitation learning), which performed best among baseline policies in simulation; ”ACT” (a policy that directly outputs joint angles); and ”Fix” (a policy outputs hard coded parameters of both action primitives) as baselines. Our method outperforms all baseline methods in terms of the final success rate which demonstrates the efficacy of our method.
Although the ACT policy shows a certain level of success in simulation, with a particularly high success rate when inserting the first endpoint, it suffers from a severe sim-to-real gap and almost never succeeds in the real world.
Fixed policy is a manually tuned policy for the test garments. It can be observed that the fixed policy still shows high success rates for inserting the left endpoint because the initial position of the clothes is relatively fixed in our experiments. However, when inserting the right endpoint, the success rate of the fixed policy drops significantly due to unforeseen deformations of the clothes during the left endpoint insertion.
In the last row of the table (”Ours(H)”), we further test the performance of our policy when the garment is initialized by a human. When the garment is initialized by the robot, the success rate shows only a slight decrease compared to human initialization, demonstrating the robustness of our policy to the initialization process.
@@ IV-E Generalization of Hangers in Real World
To further study the generalizability of our method in terms of different hangers, we report the performance of our policy when using different hangers in Table IV. Here, Hanger 2 is a hanger seen during training but without a crossbar. Hanger 3 and Hanger 1 have the same geometry, but due to difference settings in the 3D printing process, Hanger 3’s surface is much rougher. Hanger 4 is a hanger that is not seen during training.
@@
Although Hanger 2 was seen during training, there is a significant performance drop when inserting the right endpoint. As shown in Figures 5(a) and 5(b), these two images compare the state when inserting the right endpoint using different hangers. The presence of a crossbar helps the hanger separate the upper and lower layers of the clothing, making insertion easier. When using Hanger 3, as shown in Figures 5(c) and 5(d), if the hanger has too much friction, it may drag the clothes while moving, causing the clothes to become very wrinkled, which in turn reduces the success rate. When using Hanger 4, the complexity of the hanger’s shape also affects the final success rate.
Figure 5: Results with different hangers in the real world. In the real world, issues may arise when changing hangers. Figure (a) and Figure (c) show results with Hanger 1. As shown in Figure (b), when switching to Hanger 2, it becomes difficult to separate the upper and lower layers of the clothing while inserting the right end. Figure (d) illustrates that with Hanger 3, which has higher friction, the clothing tends to get wrinkled during the insertion of the left end, causing the garment to become stuck in a worse state.
V Limitation
For garments with less elasticity, such as shirts, more precise force feedback control might be required to avoid damage to the hanger and garment. Another challenge could arise when applying our action primitives to irregularly shaped garments, such as a shirt with a hood.
@@ VI Conclusion
In this work, we investigate the task of inserting a hanger into the neckline of clothes. For such a challenging but rarely studied task, we propose a simulation environment and data collection pipeline to address it. We demonstrate the efficiency of training the algorithm in a simulated environment, as well as its effectiveness and generalizability in the real world. Our pipeline facilitates efficient demonstration collecting for future robotic foundation models that require large amounts of diverse data. Future work may involve relaxing the assumptions for this task and extending the method to handle a wider variety of garments.

### v4

@@ III-F Data Collection
We observe that existing simulators, such as PyFlex [27] and MuJoCo [28], either lack the accuracy or efficiency needed for our task. The main physics accuracy issue is that they do not handle collisions and dynamics of clothing very well. Therefore, we build our simulation environment based on Taichi [11].
We use the finite element method (FEM) and implicit time integration with a single-step Newton iteration [29] for the physical simulation. Cloth self-collision and cloth-to-rigid-body collision are handled with a cubic energy function. Cloth self-intersection is prevented using Continuous Collision Detection (CCD). The collision force between cloth and rigid bodies is computed using the Signed Distance Function (SDF) of the latter. We do not perform simulation of dynamics or collision handling for rigid and articulated bodies; instead, we compute the position and velocity of objects using kinematic relationships. Depth images are rendered with SAPIEN [30], and we design the 3D models of clothes and hangers heuristically.
@@
We use an iterative data collection approach. Specifically, we first design a state-based heuristic policy to collect trajectories and use this dataset to initialize our networks. This policy achieves a success rate of approximately 50%, generating a balanced dataset with both positive and negative samples. The heuristic policy attempts to find actions that satisfy all of our predefined rules through sampling.
Next, we use the network policy to collect more data in simulation, incorporate this online data into the dataset, and iteratively fine-tune the network. During the initialization phase, we collect 12,000 trajectories. In the subsequent fine-tuning phase, we collect 1,200 trajectories per iteration, with a total of 5 iterations, resulting in 18,000 trajectories in total. Simulation and rendering take approximately 360 GPU hours on an NVIDIA RTX 3090, and training the network requires around 1 hour.
III-G Whole-body Motion Controller
@@
III-H Sim-to-real Transfer
For sim-to-real visual gap, we randomize the camera’s intrinsic and extrinsic parameters, depth maps, and masks during network training. This noise enables the network to handle a wider range of input distributions. When deploying in the real world, we use the method from [31] to obtain a refined depth map. For sim-to-real physics gap, we randomize the Young’s modulus and the density of the clothing, as well as the coefficient of friction of the garment in the simulation.
The head camera is carefully calibrated. Apart from this, we do not perform any additional real fine-tuning and directly transfer the policy learned in simulation to the real world.
@@ IV Experiment
IV-A Metrics
We evaluate the performance of the policy based on the success rate $S$ of hanger insertion. $S$ can be further divided into: the success rate of inserting the left endpoint, $S_{1}$, and the success rate of inserting the right endpoint after successfully inserting the left endpoint, $S_{2}$. In our policy, the left end is always inserted first, followed by the right end. Therefore, $S=S_{1}\times S_{2}$. The success rate is computed based on 720 trials in simulation and 16 trials in the real world.
@@ IV-B Comparison with Baselines in Simulation
We first compare our method with several state-of-the-art reinforcement learning and imitation learning approaches in simulation. Among the 144 synthetic clothing meshes, we randomly divided them into 120 training garments and 24 testing garments. The hangers used for training and testing are the ones shown in Figure 4 (a.1) and (a.2).
@@
We present the comparisons in Table I. Input refers to the type of input used by the policy, which can be either visual input (V) or state input (S). Output refers to the type of output generated by the policy, which can either be action primitives (Prim) or joint angles (Joint). Our method performs the best in simulation, achieving an overall success rate of 85.1%.
Heuristic refers to a heuristic algorithm that we manually designed based on the current mesh of the garment, as described in III-F (c). Since these rules are manually defined, their performance is not as good as methods that learn good actions from large amounts of data based on action labels.
IL stands for imitation learning, which is also based on action primitives and UNets. We treat the task as a multi-class classification task (selecting a pixel in the image) using cross-entropy loss and only retain positive samples for training the UNets. Compared with our method, IL cannot utilize failure data for training, which results in degraded performance.
SAC stands for Soft Actor-Critic algorithm [32]. In SAC, the input to the value network consists of both the observation and the action. The action is first transformed into an image where each pixel value represents the inverse of the distance to that action. This action image is then concatenated with the observation and fed into a CNN. The action network is also a CNN that directly outputs $(X_{i},Y_{i})$. Compared to our method, SAC requires learning an additional action network, while our method directly finds the maximum position on the output of the value network. The success rate is lower because the additional action network causes additional training difficulty.
ACT and DP (diffusion policy) are methods proposed by [33] and [34], respectively. We directly use data generated by heuristic policy to train an end-to-end imitation learning policy. The network’s input consists of raw depth images from the head camera, along with the masks of the clothing and the hanger. The output of the network is the angle of each robot joint. We only retain successful trajectories from the data for training, with a total of approximately 10,000 trajectories. All trajectories begin with the hanger already being grasped. Compared to our output of only 4 actions, these end-to-end policies need to output about 200 actions in a complete trajectory. These policies perform excellently when inserting the first endpoint, but due to accumulated errors and lack of data to recover from errors, they struggle with inserting the second endpoint.
@@ IV-C Ablation Studies in Simulation
w/o online data refers to training the network using only offline data (generated by the heuristic policy), without incorporating online data (generated by the UNet policy), while keeping the data quantity the same. Fine-tuning with online data results in an improvement in the final success rate. w/o neckline mask indicates training the policy without using the neckline mask. It is evident that the neckline mask also contributes to an increase in the final success rate.
@@ IV-D Comparison with Baselines in Real World
In real world experiments, we test 8 different garments and Hanger 1, as shown in Figure 4. For a fair comparison, we use the robot to initialize the garment configuration. Additionally, to evaluate our policy’s performance under different initializations, we also report the policy’s success rate with human initialization. More qualitative results are shown in Figure 5.
Figure 5: The figure shows the different garment states after applying each action primitive of RoboHanger system in the real world.
@@
We present the results of our policy compared to other methods on the real robot in Table III. We include IL (imitation learning), which performs the best among the baseline policies in simulation; ACT (a policy that directly outputs joint angles); and Fix (a policy that outputs hardcoded parameters for both action primitives) as baselines. Our method outperforms all baseline methods in terms of the final success rate, which demonstrates the efficacy of our approach.
Although the ACT policy shows a certain level of success in simulation, particularly with a high success rate when inserting the first endpoint, it suffers from a severe sim-to-real gap and rarely succeeds in the real world.
The fixed policy is a manually tuned policy designed for the test garments. We find that the fixed policy still achieves high success rates for inserting the left endpoint, as the initial position of the garment remains relatively fixed in our experiments. However, when inserting the right endpoint, the success rate of the fixed policy drops significantly due to unforeseen deformations of clothes during insertion of the left endpoint.
In the last row of the table (Ours(H)), we further report the performance of our policy when the garment is initialized by a human. When the garment is initialized by the robot, the success rate shows only a slight decrease compared to human initialization, demonstrating the robustness of our policy to variations in the initialization process.
@@ IV-E Generalization of Hangers in Real World
To further study the generalizability of our method across different hangers, we report the performance of our policy using various hangers in Table IV. Here, Hanger 2 is a hanger seen during training but without a crossbar. Hanger 3 and Hanger 1 have the same geometry, but Hanger 3 has a much rougher surface. Hanger 4 is a hanger that is not seen during training.
@@
Although Hanger 2 was seen during training, there is a significant performance drop when inserting the right endpoint. As shown in Figures 6(a) and 6(b), these two images compare the state when inserting the right endpoint using different hangers. The presence of a crossbar helps the hanger separate the upper and lower layers of the clothing, making insertion easier. When using Hanger 3, as shown in Figures 6(c) and 6(d), if the hanger has too much friction, it may drag the clothes while moving, causing the clothes to become very wrinkled, which in turn reduces the success rate. When using Hanger 4, the complexity of the hanger’s shape also affects the final success rate.
Figure 6: Results with different hangers in the real world. Issues may arise when changing hangers in reality. Figure (a) and Figure (c) show results with Hanger 1. As shown in Figure (b), when switching to Hanger 2, it becomes difficult to separate the upper and lower layers of the clothing while inserting the right end. Figure (d) illustrates that with Hanger 3, which has higher friction, the clothing tends to get wrinkled during the insertion of the left end, causing the garment to become stuck in a worse state.
V Failure Cases and Limitations
Failure Cases. For the insertion of the first endpoint, the main failure mode is the inability to separate the front and back layers of the garment. The primary reason is the complex contact forces between the garment and the hanger. For inserting the second endpoint, the main failure mode is the inability to rotate the hanger into the garment neckline. The primary reason is the large deformation of the garment at this stage, which increases the sim-to-real gap.
Limitations. Currently, garment grasping in the simulation is simplified as the garment being attached to the gripper. This physical modeling limits the exploration of potentially better strategies. A more fine-grained simulation of garment-gripper contact could further improve the overall success rate. For garments with less elasticity, such as shirts, more precise force feedback control may be required to avoid damaging both the hanger and the garment. Additionally, closed-loop control could also be introduced to help recover from failure cases.
@@ VI Conclusion
In this work, we explore the task of inserting a hanger into the neckline of an unseen garment. For this challenging task, we propose a simulation environment and a data collection pipeline. We demonstrate the efficiency of training the algorithm in a simulation environment, as well as its effectiveness and generalizability in the real world. Our pipeline enables the efficient collection of demonstrations for future robotic foundation models that require large amounts of diverse data. Future work may involve relaxing the assumptions made for this task and extending the method to handle a broader variety of garments. In addition, exploring the input of RGB images could be an interesting research direction.